24 is ____% of 40.
Answer by stanbon(75887)

24 is ____% of 40
-------------------
Equation:
24 = (x/100)40
x/100 = 24/40
--
x = 100(3/5)
x = 60
---------
Ans: 60%
===============
Cheers,
Stan H.